### v3

# A Unified Formal Theory on the Logical Limits of Symbol Grounding

Abstract
This paper synthesizes a series of formal proofs to construct a unified theory on the logical limits of the Symbol Grounding Problem. We demonstrate through a four-stage argument that meaning within a formal system must arise from a process that is external, dynamic, and non-algorithmic. First, we prove that any purely symbolic system, devoid of external connections, cannot internally establish a consistent foundation for meaning due to self-referential paradoxes. Second, we extend this limitation to systems with any finite, static set of pre-established meanings, proving they are inherently incomplete. Third, we demonstrate that the grounding process is logically incomplete; specifically, the ’act’ of connecting internal symbols to novel, emergent external meanings cannot be a product of logical inference within the system but must be an axiomatic, meta-level update. Finally, we prove that any attempt to automate this update process using a fixed, external "judgment" algorithm will inevitably construct a larger, yet equally incomplete, symbolic system. Together, these conclusions formally establish that the grounding of meaning is a necessarily open-ended, non-algorithmic process, revealing a fundamental, Gödel-style limitation for any self-contained intelligent system.
Keywords: Symbol Grounding Problem, Formal Proof, Gödel’s Incompleteness Theorems, Self-Reference, Dynamic Semantics, Embodied Cognition, Metalogic, Limitations of Computationalism.
1 Introduction
The Symbol Grounding Problem, famously articulated by Harnad[2], questions how symbols within a formal system (like a dictionary or a computer program) can acquire intrinsic meaning, rather than being endlessly defined by other ungrounded symbols. This paper aims to provide a single, unified, and formally rigorous answer by demonstrating that the problem’s solution cannot lie within any closed formal system. We will prove that grounding is not a property or state that a system can achieve and finalize, a continuous, open-ended, and non-algorithmic process.
Our argument unfolds in a logical progression over four main sections, synthesizing a series of core theorems.
First, in Section 2, we establish the foundational formal framework and prove that a purely symbolic system—one with no external connections—is logically incapable of self-grounding. By constructing a self-referential sentence about "groundability," we show that any such system with sufficient expressive power falls into inconsistency or incompleteness, akin to Gödel’s and Tarski’s limitative theorems[1, 6].
Second, in Section 3, we address the seemingly obvious solution: providing the system with an initial, finite set of grounded symbols derived from experience. We prove that any statically grounded system is also incomplete. A system can always formulate new, provable truths about its own limitations that cannot be grounded in its original, static experiential base. This establishes the necessity for grounding to be a dynamic process.
Third, in Section 4, we investigate the nature of this dynamic process. We prove that the grounding process as a whole is logically incomplete. While some routine connections might be automated, we prove that the crucial act of grounding novel, system-transcending truths is non-inferable. No general internal rule can govern this process without risking contradiction, suggesting that these key grounding acts are axiomatic, meta-level updates, not logical deductions.
Finally, in Section 5, we confront the ultimate computationalist question: can this meta-level update process be simulated by a fixed, external, algorithmic "judgment system"? We prove that any such attempt is futile. The combination of the original system and the judgment system forms a larger, but still closed, "super-system" which is itself subject to the same incompleteness. This proves that the grounding process is fundamentally non-algorithmic.
Together, these four proofs paint a complete picture: a system of meaning must, by logical necessity, be an infinitely open universe. Any attempt to draw a final boundary around it or to prescribe a fixed set of rules for its growth will fail, as it cannot account for the meaning of the boundary or the rules themselves.
2 The Formal Framework and the Impossibility of Self-Grounding
We begin by defining a formal framework to precisely discuss symbols, definitions, and grounding.
Definition 2.1 (Symbolic Language).
A symbolic language $\mathcal{L}$ is a triple $\langle S,D,G\rangle$, where:
  • $S$ is a non-empty, countable set of symbols.
  • $D$ is a Definition Function, $D:S\to\mathcal{P}(S)$, which maps each symbol $s\in S$ to its direct definitional set.
  • $G$ is the Grounding Set, $G\subseteq S$. Symbols in $G$ are postulated to be directly grounded, and for all $g\in G$, we require $D(g)=\emptyset$.
Definition 2.2 (Definitional Closure).
For a symbol $s\in S$, its Definitional Closure $C(s)$ is the smallest set satisfying: (1) $s\in C(s)$, and (2) if $s^{\prime}\in C(s)$ and $s^{\prime\prime}\in D(s^{\prime})$, then $s^{\prime\prime}\in C(s)$. $C(s)$ contains all symbols used directly or indirectly to define $s$.
Definition 2.3 (Groundability).
A symbol $s\in S$ is groundable if and only if its definitional closure intersects with the grounding set: $\text{isGroundable}(s)\iff C(s)\cap G\neq\emptyset$.
[…]
Our first investigation concerns a system that is entirely self-contained.
Definition 2.4 (Purely Symbolic System).
A symbolic language $\mathcal{L}$ is a Purely Symbolic System if its grounding set is empty, i.e., $G=\emptyset$.
To prove our main theorem, we assume the system $\mathcal{L}$ has sufficient expressive power to encompass Peano Arithmetic, allowing for Gödel numbering and self-reference[1, 5].
Axiom 2.1 (Arithmetization & Expressibility).
[…]
Axiom 2.2 (Diagonal Lemma).
For any predicate $P(x)$ in $\mathcal{L}$, there exists a sentence $\psi$ such that the system can prove $\psi\leftrightarrow P(\ulcorner\psi\urcorner)$.
Theorem 2.1 (Impossibility of Self-Grounding).
Any purely symbolic system that satisfies Axioms 2.1 and 2.2 cannot internally provide a complete and consistent determination of groundability for all its propositions.
Proof.
  1. 1.
    In a purely symbolic system, $G=\emptyset$. Thus, for any symbol $s$, $C(s)\cap G=\emptyset$. By Definition 2.3, no symbol is groundable.
  2. 2.
    Now, assume for contradiction that the system can define a predicate $Groundable(x)$ for its own propositions. Define a new predicate $Ungr(x)\equiv\neg Groundable(x)$.
  3. 3.
    By the Diagonal Lemma (Axiom 2.2), there exists a sentence $P_{\text{ungroundable}}$ such that the system proves:
    $P_{\text{ungroundable}}\leftrightarrow Ungr(\ulcorner P_{\text{ungroundable}}\urcorner)\quad\text{which is}\quad P_{\text{ungroundable}}\leftrightarrow\neg Groundable(\ulcorner P_{\text{ungroundable}}\urcorner)$
    This sentence asserts its own ungroundability.
  4. 4.
    We analyze the groundability of $P_{\text{ungroundable}}$.
    • Case A: Assume $P_{\text{ungroundable}}$ is groundable. Then $Groundable(\ulcorner P_{\text{ungroundable}}\urcorner)$ is true. The equivalence implies the system proves $\neg P_{\text{ungroundable}}$. A consistent system cannot have a true, groundable proposition whose negation it proves. This is a contradiction.
    • Case B: Assume $P_{\text{ungroundable}}$ is ungroundable. Then $\neg Groundable(\ulcorner P_{\text{ungroundable}}\urcorner)$ is true. The equivalence implies the system proves $P_{\text{ungroundable}}$. Here, the system proves a sentence whose content asserts its own lack of a semantic foundation. This creates a rupture between syntactic provability and semantic groundability, as a consistent system should not prove sentences that are, by its own definitions, meaningless.
  5. 5.
    Both cases lead to a contradiction or paradox. Therefore, a purely symbolic system cannot consistently and completely define its own groundability predicate.
This theorem establishes that meaning cannot arise from a closed loop of definitions, a conclusion that echoes the foundational crises addressed by Tarski and Kripke[6, 3]. It necessitates an external element: a non-empty grounding set $G$.
3 The Incompleteness of Statically Grounded Systems
The natural solution is to provide the system with a non-empty, but finite, grounding set $G$ based on a set of initial experiences. We call such a system "statically grounded."
Definition 3.1 (Statically Grounded System).
A symbolic language $\mathcal{L}=\langle S,D,G\rangle$ is a Statically Grounded System if its grounding set $G$ is non-empty and finite, i.e., $0<|G|<\infty$.
However, we will now prove that even this is insufficient. Any static foundation will eventually prove to be incomplete.
Theorem 3.1 (Limitation of Static Grounding).
For any consistent, statically grounded system $\mathcal{L}$, there always exists a sentence $P_{new}$ that is provable in $\mathcal{L}$, but $P_{new}$ itself cannot be grounded by the set $G$.
Proof.
  1. 1.
    Let $\mathcal{L}$ be a statically grounded system with its finite grounding set $G$.
  2. 2.
    Since $G$ is finite, the property of "being groundable by GG" is recursive. Therefore, the system can express a specific predicate $Groundable_{G}(x)$, which is true if and only if the formula with Gödel number $x$ is groundable by the set $G$. (See Appendix A).
  3. 3.
    We apply the Diagonal Lemma to the predicate $\neg Groundable_{G}(x)$. This yields a sentence $P_{new}$ such that the system proves:
    $P_{new}\leftrightarrow\neg Groundable_{G}(\ulcorner P_{new}\urcorner)$
    The informal reading of $P_{new}$ is: "This sentence is not groundable by the set G".
  4. 4.
    We analyze this sentence in the metalanguage:
    • Case A: Assume $P_{new}$ is groundable by $G$. Then $Groundable_{G}(\ulcorner P_{new}\urcorner)$ is true. The equivalence implies the system proves $\neg P_{new}$. This contradicts the system’s consistency. Thus, $P_{new}$ cannot be groundable by $G$.
    • Case B: Assume $P_{new}$ is not groundable by $G$. Then $\neg Groundable_{G}(\ulcorner P_{new}\urcorner)$ is true. The equivalence implies the system proves $P_{new}$.
  5. 5.
    From the assumption of consistency, we conclude two things: (1) The system proves $P_{new}$, and (2) $P_{new}$ is not groundable by $G$.
  6. 6.
    This completes the proof. We have constructed a sentence that is true within the system’s logic, yet its meaning transcends the system’s entire original experiential basis $G$.
This theorem reveals that grounding cannot be a one-time event. To understand the new truth $P_{new}$, the system must expand its grounding set, for example, by acquiring a new experience related to its own cognitive limitations. This implies that the grounding set $G$ must be dynamically expanding. The argument’s structure is a direct application of Gödel’s first incompleteness theorem to the problem of semantics[1, 5].
4 The Logical Incompleteness of the Grounding Act
If the grounding set must expand, what is the nature of this expansion? While some "routine" grounding acts might be automated by internal logic, is the grounding process as a whole reducible to logical deduction? We will now prove that it is not, by showing that there exists no general logical rule that can handle all grounding acts, and therefore, non-inferable grounding acts must exist.
Definition 4.1 (Deductive Symbolic System).
A deductive symbolic system $\mathcal{S}$ is a quadruple $\langle S,D,G,\vdash\rangle$, where $\vdash$ is a provability relation defined by axioms and rules of inference.
Definition 4.2 (The Grounding Act).
A Grounding Act $A(s,g)$ is a meta-level operation that updates the system’s definition function, connecting an internal symbol $s$ to a grounded symbol $g$. This act can take two forms:
  1. 1.
    G-Expansion: $g$ is a new symbol not previously in $S$ or $G$, representing a new external experience. $G$ is updated to $G^{\prime}=G\cup\{g\}$.
  2. 2.
    G-Internal Connection: $g$ is an existing symbol already in $G$. The definition function $D$ is updated, e.g., $D(s)^{\prime}=D(s)\cup\{g\}$.
Theorem 4.1 (Incompleteness of Internal Logical Grounding).
In a consistent deductive symbolic system $\mathcal{S}$, there exists no general, internal, and condition-triggered logical rule ($\vdash$) that can deduce a command to execute a grounding act for all provable but ungrounded sentences.
Proof.
We must demonstrate that no such general internal rule can exist, by analyzing both forms of the grounding act.
  1. 1.
    Case 1: Grounding via G-Expansion ($g\notin G$).
    Assume for contradiction that a general internal rule $\vdash$ can deduce the command $A(s,g)$, where $g$ is a new symbol representing an external experience not yet in $S$ or $G$.
    This is an information-theoretic contradiction. Logical inference ($\vdash$) is a closed computational process that operates only on the axioms and symbols already defined within the system $\mathcal{S}=\langle S,D,G,\vdash\rangle$.
    The symbol $g$, representing new external information, is by definition outside of this formal universe. The system’s internal rules cannot deduce a conclusion that contains a symbol ($g$) which it cannot access, name, or know. Therefore, any act that expands $G$ must be an external, axiomatic update, not an internal inference.
  2. 2.
    Case 2: Grounding via G-Internal Connection ($g\in G$).
    This is the more subtle case. Assume for contradiction that there exists a general, internal "Grounding Principle" rule in $\mathcal{S}$. This rule would state:
    "For any provable sentence $\phi$, if $\neg Groundable_{G}(\ulcorner\phi\urcorner)$ is true (i.e., $\phi$ is ungrounded), then the system must deduce ($\vdash$) a command to execute $A(\phi,g)$ for some relevant, existing $g\in G$."
    Now, let us apply this assumed general rule to the sentence $P_{new}$ from Theorem 3.1.
    • Initial State ($\mathcal{S}$): The system has proven $\mathcal{S}\vdash P_{new}$. The premise of the rule $\neg Groundable_{G}(\ulcorner P_{new}\urcorner)$ is true.
    • Action: The (assumed general) Grounding Principle rule fires. The system deduces a command and executes the grounding act $A(P_{new},g_{existing})$, where $g_{existing}$ is some symbol already in $G$. This updates the definition function $D$ such that $g_{existing}\in C(P_{new})$.
    • New State ($\mathcal{S}^{\prime}$): The system is now in state $\mathcal{S}^{\prime}$. Crucially, the grounding set itself is still G.
    • The Contradiction: We re-evaluate the predicate $Groundable_{G}(\ulcorner P_{new}\urcorner)$ in state $\mathcal{S}^{\prime}$.
      • The definitional closure $C(P_{new})$ now contains $g_{existing}$.
      • The intersection $C(P_{new})\cap G$ is no longer empty, because $g_{existing}$ is in both sets.
      • Therefore, the predicate $Groundable_{G}(\ulcorner P_{new}\urcorner)$ is now true.
    • The content of the sentence $P_{new}$ (which is $\neg Groundable_{G}(\ulcorner P_{new}\urcorner)$) has thus become false.
  3. 3.
    Conclusion of Proof: We have a formal contradiction. The system $\mathcal{S}$ used an assumed general internal rule of inference ($\vdash$) to deduce an action that directly falsified one of its own proven theorems ($P_{new}$).
    A consistent logical system cannot contain an internal rule that invalidates its own conclusions. This form of self-referential paradox is structurally analogous to the proof of Gödel’s Second Incompleteness Theorem.
    Both potential forms of a general grounding rule are thus shown to be impossible. Case 1 fails on information-theoretic grounds, and Case 2 (the general internal rule) fails on grounds of logical inconsistency.
    This does not mean that no grounding act can be deduced (a restricted, specific rule for a limited subset of symbols may well be possible). However, it proves that the grounding process as a whole is not reducible to internal logic. Because no general rule exists that can handle all cases, it follows logically that there must exist grounding acts (like for $P_{new}$) that are non-inferable from within the system.
This proves that while some grounding acts may be deductive, the grounding process in its entirety is not. The grounding of novel, system-transcending truths (like $P_{new}$) is not a conclusion reached within the system’s existing rules but is analogous to adding a new axiom from outside. This provides a formal basis for the "Aha!" moments or paradigm shifts in cognition, which are not deductions but restructurings of the axiomatic foundation itself.
5 The Incompleteness of Algorithmic Judgment
The final question remains: if the grounding act is a meta-level update, can this update process be governed by a separate, more powerful, but still algorithmic, "judgment system"?
Definition 5.1 (Cognitive Super-System).
A cognitive super-system $\mathcal{S}^{*}$ consists of:
  • An Object System $\mathcal{S}_{1}=\langle S,D,G,\vdash_{1}\rangle$, representing the agent’s knowledge.
  • A Judgment System $\mathcal{S}_{2}=\langle M,R,\vdash_{2}\rangle$, an external system with a set of fixed, algorithmic judgment rules $R$. The rules in $R$ take the state of $\mathcal{S}_{1}$ as input and output grounding commands.
Theorem 5.1 (Limitation of Logical Judgment Systems).
Any cognitive super-system $\mathcal{S}^{*}$ with a fixed, algorithmic judgment system is itself an incomplete deductive system.
Proof.
  1. 1.
    The core insight is that because the judgment rules $R$ of $\mathcal{S}_{2}$ are fixed and algorithmic, the behavior of the entire super-system $\mathcal{S}^{*}$ (including all its future state updates) is completely determined by a fixed set of initial axioms and rules.
  2. 2.
    Therefore, the super-system $\mathcal{S}^{*}$ is itself a larger, more complex, but ultimately static deductive symbolic system.
  3. 3.
    We can now apply Theorem 3.1 (the limitation of static grounding) directly to this super-system $\mathcal{S}^{*}$.
  4. 4.
    For the system $\mathcal{S}^{*}$, we can construct a Gödel-style sentence $P^{*}$ with the content: "This sentence cannot be grounded by the rules of the super-system $\mathcal{S}^{*}$."
  5. 5.
    As per the proof of Theorem 3.1, we know that if $\mathcal{S}^{*}$ is consistent, then (a) $\mathcal{S}^{*}\vdash P^{*}$ (it is provable within the super-system’s logic) and (b) $P^{*}$ is not groundable by the rules of $\mathcal{S}^{*}$ (which includes the judgment rules of $\mathcal{S}_{2}$).
  6. 6.
    This means the judgment system $\mathcal{S}_{2}$, which was designed to handle all necessary grounding acts, has failed. It is faced with a truth $P^{*}$ that its own rules cannot ground. To ground $P^{*}$, one would need a yet higher-order judgment system $\mathcal{S}_{3}$, leading to an infinite regress.
  7. 7.
    This hierarchical paradox is analogous to arguments against strong AI, such as Searle’s Chinese Room, where adding more layers of symbolic processing fails to bridge the gap to semantics[4]. Therefore, any attempt to fully simulate the open, dynamic grounding process with a fixed, algorithmic system is logically bound to fail.
6 Conclusion
This paper has constructed a unified, four-part formal proof demonstrating the fundamental logical limits of symbol grounding. Our chain of reasoning leads to a single, powerful conclusion: a system of meaning must, by logical necessity, be an infinitely open universe.
We have shown that:
  1. 1.
    Meaning cannot be generated internally within a closed symbolic loop (impossibility of self-grounding). It requires an external connection.
  2. 2.
    This external connection cannot be a one-time, fixed set of experiences. Any static foundation is inherently incomplete, necessitating a dynamic process of continuous grounding.
  3. 3.
    The act of expanding this connection for novel, emergent truths is not a logical deduction. This proves that the grounding process is logically incomplete, and that non-inferable, axiomatic updates are a necessary part of it.
  4. 4.
    This update process cannot be fully captured by any fixed, external algorithm. The meta-system responsible for grounding must itself be open and non-algorithmic.
This result places a Gödelian limit on strong computationalism[1, 4], suggesting that core aspects of intelligence, such as understanding and the "Aha!" moment of insight, are not reducible to a fixed set of computations. The essence of meaning lies not in the symbols themselves or even in a static map to the world, but in the open-ended, non-algorithmic process of continuously and creatively updating that map.
Appendix A Formal Construction of the Groundability Predicate
This appendix provides a more detailed formal construction of the $Groundable(n)$ predicate within a system containing Peano Arithmetic, demonstrating that Axiom 2.1 is well-founded.
  1. 1.
    Arithmetization of Syntax: We assume standard Gödel numbering. The definition function $D$ is assumed to be recursive, so the relation $m\in D(n)$" can be represented by a recursive arithmetic predicate $Def(\ulcorner m\urcorner,\ulcorner n\urcorner)$. The predicate for membership in the grounding set, $is\_in\_G(\ulcorner g\urcorner)$, is syntactically well-formed, even if its extension is empty (as in a purely symbolic system).
  2. 2.
    Arithmetization of Definitional Closure: The definitional closure $C(s)$ is the transitive closure of the relation defined by $D$. The transitive closure of a recursive relation is itself recursive. We can define a predicate $InClosure(m,n)$ which is true iff $\ulcorner m\urcorner\in C(\ulcorner n\urcorner)$. This is expressible by stating that there exists a finite sequence (coded as a single number) $s_{0},s_{1},\dots,s_{k}$ such that $\ulcorner s_{0}\urcorner=n$, $\ulcorner s_{k}\urcorner=m$, and for all $i<k$, $Def(\ulcorner s_{i+1}\urcorner,\ulcorner s_{i}\urcorner)$ holds.
  3. 3.
    Definition of Groundable(n): With the above components, we can formally define the $Groundable(n)$ predicate as:
    $Groundable(n)\equiv\exists g(is\_in\_G(g)\land InClosure(g,n))$
    Since all components of this formula are recursive, the predicate $Groundable(n)$ is itself recursive and thus expressible within the system.
Appendix B Glossary of Key Terms
Term Definition
Symbolic Language A formal system consisting of symbols and definition rules.
Grounding Set ($G$) The set of symbols whose meanings are given externally.
Purely Symbolic System A system where the grounding set is empty ($G=\emptyset$).
Definitional Closure ($C(s)$) The set of all symbols used directly or indirectly to define a symbol $s$.
Groundability The property of a symbol whose definitional closure intersects with the grounding set.
Semantic Deferral The state of an ungroundable symbol whose meaning is endlessly deferred.
Diagonal Lemma A theorem that allows a formal system to construct self-referential sentences.